OpenStudy (anonymous):

please solve x + y = –1 y = 4 x

OpenStudy (lgbasallote):

if y = 4x then just substitute x + y = -1 x + 4x = -1 5x = -1 solve for x

OpenStudy (anonymous):

well see this is the thing, it has to be graphed and thats the tricky part

OpenStudy (lgbasallote):

then graph x + y = -1 then graph y = 4x the intersection is the value of x and y

OpenStudy (anonymous):

no is there many solutions are no solutions